# Spacetime Gaussian Feature Splatting for Real-Time Dynamic View Synthesis

### 28 Dec 2023 v1

@@ Appendix D More Results
We provide per-scene quantitative comparisons on the Neural 3D Video Dataset [45] (Tab. 7), Google Immersive Dataset [10] (Tab. 8) and Technicolor Dataset [71] (Tab. 9). Our method outperforms the other baselines on most scenes.
@@
Figs. 7, 8 and 9 show more visual comparisons of our full model and our lite-version model with NeRFPlayer [75], HyperReel [4], K-Planes [26], MixVoxels-L [81] and Dynamic 3DGS [59] on the Neural 3D Video Dataset [45].
@@
Appendix E Experiment Details
Since the open source code of MixVoxels [81] does not contain the training config for MixVoxels-X, we use MixVoxels-L in our comparisons. We train HyperReel with their official code to generate visual examples. Note that the training time of HyperReel for each scene on the Technicolor Dataset is about 3.5 hours, while that of our full model on the Technicolor Dataset is only about 1 hour. For Dynamic 3DGS, its performance on the Neural 3D Video Dataset is not reported in their original paper. When applying its open source code to Neural 3D Video Dataset with their default hyperparameters, the rendering quality is subpar. So we tune its hyperparameters to improve the performance on this dataset. We use Nvidia RTX 3090 when reporting our rendering speed in the comparisons with other methods, and use Nvidia RTX 4090 in our real-time demos.
We use the original centered-undistorted camera model from 3DGS [39] for the Neural 3D Video Dataset. We implement the uncentered-undistorted camera model for the Technicolor Dataset. For the Google Immersive Dataset [10], to evaluate on the same distorted videos as [2, 75], we further adapt our method to fit the uncentered-distorted camera model with a differentiable image space warping, which maps perspective view to fish-eye distorted view. For the real-time demo, we retrain our models on the undistorted videos for simplicity.
For the Theater, Train and Birthday scenes in Tabs. 3 and 9, we only use 25 percent of SfM points from the frames (except the first frame whose SfM points are all used). In the ablation study on the number of frames whose SfM points are used (Tab. 5), we use all the points in each sampled frames.

### 04 Apr 2024 v2

@@ Appendix F More Results
We provide per-scene quantitative comparisons on the Neural 3D Video Dataset [48] (Tab. 11), Google Immersive Dataset [10] (Tab. 12) and Technicolor Dataset [76] (Tab. 13). Our method outperforms the other baselines on most scenes. We also provide per-scene Gaussian numbers of trained 50-frame models in Tab. 14.
@@
Figs. 10, 11 and 12 show more visual comparisons of our full model and our lite-version model with NeRFPlayer [80], HyperReel [4], K-Planes [28], MixVoxels-L [87] and Dynamic 3DGS [64] on the Neural 3D Video Dataset [48].
@@
Table 14: Per-scene Gaussian numbers (K) on three datasets. For each scene, the number is averaged over 50-frame models.
Avg. 01_Welder 02_Flames 04_Truck 09_Exhibit 10_Face_Paint_1 11_Face_Paint_2 12_Cave
Google Immersive Dataset [10] 427 571 389 374 484 285 249 629
Avg. Coffee Martini Cook Spinach Cut Roasted Beef Flame Salmon Flame Steak Sear Steak
Neural 3D Video Dataset [48] 215 262 189 169 319 177 176
Avg. Birthday Fabien Painter Theater Trains
Technicolor Dataset [76] 374 379 295 412 313 470